How to calc. value of (25.364*61.65)/.091 using log tables? Explain in detail.

Grade:12

1 Answers

noel thekkekara
12 years ago

logs are to the base 10 ,as 25.364 is between 10^1 to 10^2 1.-----. then find the log for 0.2536 since log can be used only for 4 nos.

61.65 is between 10^1 to 10^2 1.-----. then find the log for 0.6165. thereafter add both their logs

log(a*b)=log a+log b

(1.4031+10)+(1.7896+4)

=1.4041+1.7900

=3.1941

and find the antilog of 3.1941

3194

antilog=0.1563 ,as 3 was present

0.1563*10*10*10*10=1563

now you want 1563/0.91

1563 betw 10^3 & 10^4

log is 3.1941

write 0.091 =0.91/10

mutiply 10in the numerator

4.1941-0.9590=3.2351

antilog=2351's antilog=0.1718*10*10*10*10*10*10

your ans is approx 17180